OpenStudy (anonymous):

A contractor employed 210 men to build a house in 60 days. After 12 days , he was joined by 70 more men. In how many days will the remaining work be finished.

OpenStudy (chmvijay):

for 210 men it takes 60 days for 210+70 men it takes = 280*60 /210

OpenStudy (ranga):

After 12 days, if the remaining work were continued with the original 210 men the job would have finished in (60 - 12 = 48) days. But with 280 men the job will be finished in 48 * 210/280 = 36 days.

OpenStudy (wolf1728):

210 men build a house in 60 days 210 men in 1 day build 1/60 house After 12 days, 12/60 or 1/5 of the house is built. Then 280 men work to finish 4/5 of a house 280 men can finish a house in 210/280 * 60 = 3/4 * 60 = 45 days They can finish 4/5 of a house in 45 * 4/5 = 36 days So total days worked = 12 + 36 = 48 days

hero (hero):

The total days worked is 48 days. We can all agree to that, but as far as answering the actual question "In how many days will the remaining work be finished?". 36 days is correct. You start with 210 men and it takes the 60 days to do the job so that's 210*60 = 12600 units of work to be done But they only worked for 12 days so 210*12 = 2520 units of work completed so far Which means 12600 - 2520 = 10080 units of work remain If 70 more men join, that's 280 men working now. If we divide 10080/280 we will get how many days it takes to complete the remainder of the work 10080/280 = 36 Therefore, it takes 36 days to complete the remainder of the work.

OpenStudy (wolf1728):

Agreed that it is 36 days (That is the next to the last line in my answer. Guess I should have made it the last line. LOL)

hero (hero):

Yes, usually the last thing you post is the conclusion sentence where you state the actual answer to the question.